Solve the integral equation: y(t)1- y(t-)e2 dr. y(0) 1
Answer #1

Here the given integral equation is y (t) = 1- Syet-e).se 27 dt with 101-1 Now t Sye-T)eude= y(t) *e r By convolution properPFI Pal = pel + Bari ... LfYx) = 3 - PA PA taking in verse Laplace transformation y(t) = L-157] -2-15 pt) = 2-et. y(t) = 2-et